OpenStudy (anonymous):

Simple Interest I = $15 r = 8% t = 3 quarters p = ?

OpenStudy (akashdeepdeb):

PRT/100 = SI So P = S1 * 100/RT Understood? :)

OpenStudy (anonymous):

not really.. cant i just use the formula I=Prt ?

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (akashdeepdeb):

Well you just shift the terms around you know like this! If ab = z and we have to find b Then b = z/a Like that! :D

OpenStudy (anonymous):

i get it but what about the t ? its suppose to be num of years right? so do i calculate t=3/4 since it gives t=3 quarters @AkashdeepDeb

OpenStudy (akashdeepdeb):

Yes use it as 3 quarters or 3/4 Because 1 year has 3months 4 times and 1 quarter in a year is 3/12 So 3 quarters is 9/12 which 3/4 itself!! :D Understood? :)

OpenStudy (anonymous):

yeah thanks! so lets say they give t=30 days so i should calculate it as t=30/365 right?

OpenStudy (akashdeepdeb):

EXCELLENT!! :D